OpenStudy (anonymous):

lim x→-2 x+2/√(x^2-4)

OpenStudy (anonymous):

@Loser66 what's the l'hopital rule?

OpenStudy (anonymous):

do i multiply both of them by √(x^2 +2)

OpenStudy (anonymous):

@Loser66 i don't know what derivative means

OpenStudy (loser66):

oop, you study limit and not know about derivative? ok, I take it back and think other way, hihi, sorry for that

OpenStudy (anonymous):

@Loser66 i don't pay attention in class

OpenStudy (loser66):

sorry friend, I really don't know how to solve without applying l'hopital rule. wait for other 's help,

OpenStudy (anonymous):

\[\begin{align*}\lim_{x\to-2}\frac{x+2}{\sqrt{x^2-4}}&=\lim_{x\to-2}\frac{x+2}{\sqrt{(x+2)(x-2)}}\\ &=\lim_{x\to-2}\frac{x+2}{\sqrt{x+2}\sqrt{x-2}}\\ &=\lim_{x\to-2}\frac{\sqrt{x+2}}{\sqrt{x-2}}\\ \end{align*}\]

OpenStudy (anonymous):

@SithsAndGiggles would i plug in the -2 for the x next

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

@SithsAndGiggles thanks

OpenStudy (anonymous):

You're welcome